A ramp that is 18 meters long makes an angle of 9º with the ground. What is the height h of the ramp? Round your answer to the nearest tenth of a meter.
Answer by jim_thompson5910(35256)

sin(theta) = opposite/hypotenuse

sin(9) = h/18

18*sin(9) = h

h = 18*sin(9)

h = 18*0.156434465040231

h = 2.81582037072416


So the height of the ramp is roughly 2.81582037072416 m which rounds to 2.8 meters
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------